If the K a of a monoprotic weak acid is 1.4 × 10 − 6 , what is the pH of a 0.11 M solution of this acid
Answer #1

Assume a weak monoprotic acid is HA. Consider dissociation of weak acid HA in water.

HA (aq) + H2O (l) H3O + (aq) + A - (aq)  

Equilibrium constant for above reaction is, K a = [ H3O + ] [ A - ] / [ HA ] = 1.4 10 -06

Let's use ICE table.

Concentration

(M)

HA (aq) H3O + (aq) + A - (aq)  
I 0.11
C - X +X +X
E 0.11- X X X

Substituting these values in K a = [ H3O + ] [ A - ] / [ HA ] = 1.4 10 -06 , we get

( X ) ( X ) / 0.11 - X =1.4 10 -06

X 2 / 0.11 - X = 1.4 10 -06  

X 2 = ( 0.11 - X ) 1.4 10 -06  

X 2 =1.54 10 -07 -  1.4 10 -06  X

X 2 +  1.4 10 -06  X - 1.54 10 -07 = 0

Comparing above equation with a X 2 + b X+ c = 0 , we get a = 1 , b =1.4 10 -06 , c = - 1.54 10 -07

Now , solve equation for X.

X = - b +/- b 2 -4 ac / 2 a  

X = - 1.4 10 -06 + /- ( 1.4 10 -06 ) 2 - 4 (1) ( - 1.54 10 -07 ) / 2 (1)

X = - 1.4 10 -06 + /- 6.16 10 -07 / 2

X = - 1.4 10 -06 + /- 7.85 10 -04 / 2

X = - 1.4 10 -06 + 7.85 10 -04 / 2 = 0.000392 M

Or X = - 1.4 10 -06 - 7.85 10 -04 / 2 = - 0.000393 M

We can accept X = 0.000392 M

We have, X = [ H3O + ] = [ A - ] = 0.000392 M

We have relation, pH = - log [ H3O + ]

pH = - log ( 0.000392 )

pH = 3.41